# Scalable and robust regression models for continuous proportional data

### v1

@@ Acknowledgement
The MMI data analysis involves an NNGP prior for the spatial random effects. To enable nearest neighbor calculation using the spNNGP R package (Finley et al.,, 2022), we used WGS 84 / UTM zone 15N (EPSG:32615) coordinate system. The NNGP prior requires the specification of an ordering and a number of neighborhoods. Following the default setting of spNNGP R package, we use coordinate-based ordering and use 15 nearest neighbors. We use an exponential covariance kernel $\operatorname*{cov}\{u(s),u(s^{\prime})\}=\sigma_{u}^{2}\exp(-\|s-s^{\prime}\|/\rho)$ in the construction of the NNGP prior. We fixed the spatial range parameter at $\rho=200$km, leading to an effective range of approximately 600km, the distance where the spatial correlation is below 0.05 (before NNGP approximation). This specification is consistent with the analysis of Fox et al., (2020), where the empirical spatial autocorrelation of stream benthic macroinvertebrate MMI in the US was found to be close to zero for distances beyond 580km.
In the comparative analysis shown in Table 3 in the main manuscript, the lake with a zero MMI (COMID: 22721231) was excluded since beta and cobin regressions do not accommodate an exact 0 response. Table S.4 summarizes the spatial micobin regression results using the full dataset ($n=950$), along with results from Table 3 ($n=949$) for a comparison. The estimates and credible intervals are highly similar to each other. Furthermore, the quantile residual plot in Figure 4 indicates that two low-MI lakes (COMID: 9201925, 22845861) are highly influential in the beta and cobin regression fit with $n=949$. After removing these two lakes, the beta, cobin, micobin regression results based on the data with $n=947$ are presented in Table S.5.
For the beta regression results, the selected variables based on 95% credible intervals not including zero changed; bfi is no longer selected, while agkffact and conif are now selected. For the cobin and micobin regressions, there were no changes in selected variables. In terms of the difference in the posterior mean estimate $\hat{\bm{\beta}}$ compared to the result with data of size $n=949$, measured by the Euclidean distance, the beta regression exhibits the largest change, with a value of 0.743 (0.516 excluding the intercept). The cobin regression shows a change of 0.444 (0.273 excluding the intercept), and the micobin regression is the most stable, with a change of 0.069 (0.040 excluding the intercept).
Table S.4: MMI data analysis results with original data with size $n=950$ using micobin regression, along with the result from Table 3 ($n=949$) for a comparison. Results where 95% credible intervals that do not include zero are highlighted in bold.
Micobin regression $(n=950)$ Micobin regression $(n=949)$
Variable Estimate 95% CI Estimate 95% CI
(Intercept) -1.758 (-3.517, -0.04) -1.797 (-3.551, -0.085)
agkffact -3.456 (-6.175, -0.794) -3.457 (-6.113, -0.800)
bfi 0.219 (-0.100, 0.537) 0.229 (-0.082, 0.548)
cbnf 0.187 (-0.040, 0.415) 0.191 (-0.035, 0.425)
conif 0.128 (0.048, 0.208) 0.123 (0.044, 0.203)
crophay -0.060 (-0.222, 0.101) -0.054 (-0.213, 0.105)
fert -0.071 (-0.296, 0.13) -0.082 (-0.300, 0.138)
manure -0.031 (-0.178, 0.115) -0.029 (-0.173, 0.118)
pestic97 -0.023 (-0.106, 0.059) -0.025 (-0.108, 0.059)
urbmdhi -0.141 (-0.243, -0.038) -0.142 (-0.243, -0.041)
@@
agkffact, soil erodibility factor; bfi, base flow index; cbnf, cultivated biological N fixation; conif, coniferous forest cover; crophay, crop/hay land cover; fert, synthetic N fertilizer use; manure, manure application; pestic97, 1997 pesticide use; urbmdhi, medium/high-density urban land cover. All variables are $\log_{2}(x+1)$ transformed.
Table S.5: Comparison of MMI data analysis results with three different models with dataset of size $n=947$. Variables where 95% credible intervals that do not include zero are highlighted in bold. The last row corresponds to the Euclidean distance between the estimates based on $n=949$ and $n=947$.
$(n=947)$ Beta regression Cobin regression Micobin regression
Variable Estimate 95% CI Estimate 95% CI Estimate 95% CI
(Intercept) -1.829 (-3.621, -0.070) -1.756 (-3.531, 0.006) -1.741 (-3.520, 0.018)
agkffact -3.088 (-5.982, -0.206) -3.150 (-6.019, -0.258) -3.494 (-6.220, -0.822)
bfi 0.244 (-0.075, 0.568) 0.228 (-0.088, 0.552) 0.219 (-0.097, 0.540)
cbnf 0.191 (-0.044, 0.430) 0.200 (-0.035, 0.437) 0.196 (–0.034, 0.424)
conif 0.096 (0.014, 0.175) 0.103 (0.021, 0.183) 0.125 (0.045, 0.204)
crophay -0.057 (-0.223, 0.110) -0.053 (-0.218, 0.114) -0.050 (-0.210, 0.110)
fert -0.096 (-0.327, 0.135) -0.104 (-0.329, 0.122) -0.089 (-0.316, 0.135)
manure -0.001 (-0.148, 0.148) -0.009 (-0.157, 0.138) -0.022 (-0.167, 0.122)
pestic97 -0.031 (-0.118, 0.057) -0.030 (-0.119, 0.057) -0.027 (-0.110, 0.055)
urbmdhi -0.180 (-0.283, -0.076) -0.169 (-0.275, -0.064) -0.143 (-0.242, -0.043)
Change $\|\hat{\bm{\beta}}^{(949)}-\hat{\bm{\beta}}^{(947)}\|_{2}=0.743$ $\|\hat{\bm{\beta}}^{(949)}-\hat{\bm{\beta}}^{(947)}\|_{2}=0.444$ $\|\hat{\bm{\beta}}^{(949)}-\hat{\bm{\beta}}^{(947)}\|_{2}=0.069$
agkffact, soil erodibility factor; bfi, base flow index; cbnf, cultivated biological N fixation; conif, coniferous forest cover; crophay, crop/hay land cover; fert, synthetic N fertilizer use; manure, manure application; pestic97, 1997 pesticide use; urbmdhi, medium/high-density urban land cover. All variables are $\log_{2}(x+1)$ transformed.

### v2

@@ Acknowledgement
The MMI data analysis involves an NNGP prior for the spatial random effects. To enable nearest neighbor calculation using the spNNGP R package (Finley et al., , 2022), we used the WGS 84 / UTM zone 15N (EPSG:32615) coordinate system. The NNGP prior requires the specification of an ordering and a number of neighborhoods. Following the default setting of spNNGP R package, we use coordinate-based ordering and use 15 nearest neighbors. We use an exponential covariance kernel $\operatorname*{cov}\{u(s),u(s^{\prime})\}=\sigma_{u}^{2}\exp(-\|s-s^{\prime}\|/\rho)$ in the construction of the NNGP prior. We fixed the spatial range parameter at $\rho=200$km, leading to an effective range of approximately 600km, the distance where the spatial correlation is below 0.05 (before NNGP approximation). This specification is consistent with the analysis of Fox et al., (2020), where the empirical spatial autocorrelation of stream benthic macroinvertebrate MMI in the US was found to be close to zero for distances beyond 580km.
S.6.2 Quantile residual analysis for identifying potential outliers
Figure S.7: Comparison of quantile residuals for goodness-of-fit assessment, along with two observations corresponding to the lowest quantiles. The red line corresponds to the $y=x$ line.
We first present the quantile residual (Dunn and Smyth, , 1996) analysis results from four different models based on the dataset $\mathcal{D}^{\circ}$, where quantile residuals are computed conditional on the spatial random effects, with model parameters fixed at the posterior mean (posterior median of $\lambda$ for cobin). Figure S.7 shows that beta and cobin regression exhibit a lack of fit in the left tail, whereas beta rectangular and micobin regression capture the left tail accurately but show slightly degraded fit in the right tail.
Further inspection reveals that the two observations with the lowest quantile residuals are identical for all four models, with MMI values of 0.02 and 0.021. This suggests that these lakes are not well explained by the beta and cobin models, as their low MMI values fall outside the expected range of residual variation under those models. In contrast, beta rectangular and micobin regression accommodate these low values more naturally, but at the same time imply greater dispersion in the right tail than what is observed in the data. This reflects the fundamental characteristic of the beta rectangular and micobin models based on mixture formulation, which allows for improved robustness to extreme values, but at the cost of less sharp probabilistic predictions. See also Figure S.6 for similar behavior of quantile residuals.
In summary, the quantile residual analysis identifies two low MMI lakes as potential outliers under the beta and cobin regression models. One of the key goals of the MMI data analysis is to evaluate and compare the robustness of the models in the presence of such outliers.
S.6.3 Additional summary and figures
Table S.7: Estimated fixed-effect coefficients $(\hat{\bm{\beta}}^{\star})$ from two spatial regression models (beta rectangular, micobin) fitted on $\mathcal{D}^{\star}(n=950)$ and 95% credible intervals. Bold entries indicate non-intercept coefficients whose 95% credible intervals exclude zero.
Betarec regression ($\|\hat{\bm{\beta}}^{\star}-\hat{\bm{\beta}}\|_{2}=0.103$ ) Micobin regression ($\|\hat{\bm{\beta}}^{\star}-\hat{\bm{\beta}}\|_{2}=0.048$ )
Variable $\hat{\beta}_{j}^{\star}$ ($n=950$) $\hat{\beta}_{j}$ ($n=947$) $\hat{\beta}_{j}^{\star}$ ($n=950$) $\hat{\beta}_{j}$ ($n=947$)
(Intercept) -1.596 (-3.385, 0.156) -1.696 (-3.543, 0.092) -1.758 (-3.517, -0.040) -1.741 (-3.520, 0.018)
agkffact -3.072 (-5.916, -0.288) -3.082 (-5.961, -0.200) -3.456 (-6.175, -0.794) -3.494 (-6.220, -0.822)
bfi 0.202 (-0.113, 0.530) 0.220 (-0.101, 0.551) 0.219 (-0.100, 0.537) 0.219 (-0.097, 0.540)
cbnf 0.200 (-0.033, 0.437) 0.200 (-0.036, 0.437) 0.187 (-0.040, 0.415) 0.197 (-0.034, 0.424)
conif 0.107 (0.026, 0.186) 0.103 (0.023, 0.184) 0.128 (0.048, 0.208) 0.125 (0.045, 0.204)
crophay -0.036 (-0.202, 0.132) -0.042 (-0.209, 0.124) -0.060 (-0.222, 0.101) -0.050 (-0.210, 0.110)
fert -0.127 (-0.357, 0.100) -0.120 (-0.348, 0.105) -0.071 (-0.296, 0.130) -0.089 (-0.316, 0.135)
manure -0.018 (-0.166, 0.130) -0.013 (-0.165, 0.137) -0.031 (-0.178, 0.115) -0.022 (-0.167, 0.122)
pestic97 -0.034 (-0.121, 0.053) -0.034 (-0.121, 0.054) -0.023 (-0.106, 0.059) -0.027 (-0.110, 0.055)
urbmdhi -0.162 (-0.266, -0.058) -0.172 (-0.278, -0.067) -0.141 (-0.243, -0.038) -0.143 (-0.242, -0.043)
[…]
Beta Cobin Beta rectangular Micobin
(Intercept) -1.829 (-3.621, -0.070) -1.756 (-3.531, 0.006) -1.696 (-3.543, 0.092) -1.741 (-3.520, 0.018)
agkffact -3.088 (-5.982, -0.206) -3.150 (-6.019, 0.258) -3.082 (-5.961, -0.200) -3.494 (-6.220, -0.822)
bfi 0.244 (-0.075, 0.568) 0.228 (-0.088, 0.552) 0.220 (-0.101, 0.551) 0.219 (-0.097, 0.540)
cbnf 0.191 (-0.044, 0.430) 0.200 (-0.035, 0.437) 0.200 (-0.036, 0.437) 0.197 (-0.034, 0.424)
conif 0.096 (0.014, 0.175) 0.103 (0.021, 0.183) 0.103 (0.023, 0.184) 0.125 (0.045, 0.204)
crophay -0.057 (-0.223, 0.110) -0.053 (-0.218, 0.114) -0.042 (-0.209, 0.124) -0.050 (-0.210, 0.110)
fert -0.096 (-0.327, 0.135) -0.104 (-0.329, 0.122) -0.120 (-0.348, 0.105) -0.089 (-0.316, 0.135)
manure -0.001 (-0.148, 0.148) -0.009 (-0.157, 0.138) -0.013 (-0.165, 0.137) -0.022 (-0.167, 0.122)
pestic97 -0.031 (-0.118, 0.057) -0.030 (-0.119, 0.057) -0.034 (-0.121, 0.054) -0.027 (-0.110, 0.055)
urbmdhi -0.180 (-0.283, -0.076) -0.169 (-0.275, 0.064) -0.172 (-0.278, -0.067) -0.143 (-0.242, -0.043)
[…]
Table S.9: Estimated average slopes (average marginal effects) from four spatial regression models fitted on $\mathcal{D}^{\circ}(n=949)$ and 95% credible intervals. Bold entries indicate the average slopes whose 95% credible intervals exclude zero.
Beta regression Cobin regression Betarec regression Micobin regression
agkffact -0.196 (-0.422, 0.025) -0.219 (-0.433, -0.000) -0.231 (-0.449, -0.015) -0.262 (-0.463, -0.061)
bfi 0.026 (0.001, 0.051) 0.022 (-0.002, 0.046) 0.016 (-0.008, 0.040) 0.017 (-0.006, 0.041)
cbnf 0.012 (-0.006, 0.031) 0.014 (-0.004, 0.032) 0.015 (-0.003, 0.033) 0.014 (-0.003, 0.032)
conif 0.006 (-0.000, 0.012) 0.007 (0.001, 0.013) 0.008 (0.002, 0.014) 0.009 (0.003, 0.015)
crophay -0.006 (-0.019, 0.007) -0.005 (-0.018, 0.008) -0.003 (-0.016, 0.010) -0.004 (-0.016, 0.008)
fert -0.006 (-0.023, 0.012) -0.007 (-0.024, 0.010) -0.010 (-0.027, 0.007) -0.006 (-0.023, 0.010)
manure -0.004 (-0.015, 0.008) -0.003 (-0.014, 0.009) -0.001 (-0.013, 0.010) -0.002 (-0.013, 0.009)
pestic97 -0.001 (-0.008, 0.006) -0.002 (-0.008, 0.005) -0.003 (-0.009, 0.004) -0.002 (-0.008, 0.004)
urbmdhi -0.014 (-0.022, -0.006) -0.013 (-0.021, -0.005) -0.013 (-0.021, -0.005) -0.011 (-0.018, -0.003)
@@
agkffact, soil erodibility factor; bfi, base flow index; cbnf, cultivated biological N fixation; conif, coniferous forest cover; crophay, crop/hay land cover; fert, synthetic N fertilizer use; manure, manure application; pestic97, 1997 pesticide use; urbmdhi, medium/high-density urban land cover.
Tables S.7 and S.8 summarize results based on the full dataset $\mathcal{D}^{\star}$ ($n=950$) and the reduced dataset $\mathcal{D}$ ($n=947$), respectively. Based on the full dataset results in Table S.7, compared to the beta rectangular regression model, the micobin regression model exhibits greater stability of regression coefficients, improved predictive performance in terms of PSIS-LOO and WAIC on full dataset, and substantial computational advantages.
Table S.9 reports results based on the dataset $\mathcal{D}^{\circ}$ ($n=949$) in terms of average slopes, also known as average marginal effects (Williams, , 2012; Arel-Bundock et al., , 2024). These quantities are computed as the empirical average of the partial derivative $\frac{\partial E(y\mid\bm{x})}{\partial x_{j}}\big|_{\bm{x}=\bm{x}_{i}}$ across the observed data points $i=1,\dots,n$ for a given covariate $x_{j}$. Average slopes are often interpreted as an approximation of an average change of $E(y\mid\bm{x})$ associated with a unit change of $x_{j}$, holding other covariates constant (Norton et al., , 2019). As an example, the estimated average marginal effect of urbmdhi from the micobin regression is -0.011. This can be interpreted as doubling the percentage of medium/high-density urban land cover is associated with a decrease of MMI by 1.1 on average (on a 0-100 scale), holding the other covariates constant.
Figure S.8 compares spatial MMI predictions at 55,215 unsampled lakes from each model against those from the spatial micobin regression. The results show that the predictions of the spatial beta regression model are more influenced by the two lakes with low MMI, leading to a substantial underestimation of MMI of almost 14 units (on the original scale of 0-100), compared to the beta-rectangular and micobin. In contrast, the spatial cobin regression model is noticeably less influenced by these two lakes. Figure S.9 shows trace plots of log-likelihood conditional on random effects from four models based on $\mathcal{D}^{\circ}$, which form the basis for computing PSIS-LOO and WAIC.
Figure S.8: (Left) Predicted MMI at 55,215 lakes from the spatial beta (left top), spatial cobin (left middle), and spatial beta rectangular (left bottom) regression model based on $\mathcal{D}^{\circ}$ in terms of posterior predictive mean of $E\{Y(s^{*})\mid X(s^{*})\}$ at unsampled location $s^{*}$. (Right) Differences in predicted MMI between spatial micobin regression and each of the three models on the left, respectively. See Figure S.7 for the legend corresponding to triangle and square markers, indicating the location of two low MMI lakes.
Figure S.9: Trace plots of the log-likelihood (conditional on spatial random effects) from the spatial beta, cobin, beta rectangular, and micobin regression models, based on dataset $\mathcal{D}^{\circ}$. Each panel shows results from three MCMC chains.
S.7 Data augmentation and boundary-concentrated data
S.7.1 Bayesian cobin regression under boundary-concentrated settings
The Kolmogorov-Gamma augmentation scheme proposed in this article shares many similarities with the Pólya-Gamma augmentation scheme (Polson et al., , 2013), including its limitations. One important limitation of the Pólya-Gamma augmentation for logistic models arises in highly unbalanced binary data settings, where the corresponding Markov chain often exhibits poor mixing. This issue is formally investigated in Johndrow et al., (2019), where an “infinitely imbalanced” data setting (Owen, , 2007) is considered, defined as $(y_{1},y_{2},\dots,y_{n})\in\{0,1\}^{n}$ such that $\sum_{i=1}^{n}y_{i}=1$ (1 success, n1n-1 failure) with growing $n$. In this setting, Johndrow et al., (2019) considered an intercept-only logistic model $\mathrm{logit}\{\operatorname*{pr}(y_{i}=1)\}=\beta$ for $i=1,\dots,n$, and investigated the Markov chain arising from Polya-Gamma augmentation, targeting the posterior of $\beta$.
We illustrate through an example that the Markov chain based on the Kolmogorov-Gamma augmentation may exhibit poor mixing behavior when most of the data are concentrated on either 0 or 1. Without loss of generality, analogous to “infinitely imbalanced” binary data setting, we consider “infinitely boundary-concentrated” setting, defined as $(y_{1},y_{2},\dots,y_{n})\in[0,1]^{n}$ satisfying $\sum_{i=1}^{n}y_{i}=1$, so that the empirical average of response is $1/n$ with $n$ datapoints. In this setting, we consider the intercept-only cobin regression model $Y_{i}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}\mathrm{cobin}(\beta,\lambda^{-1})$ with cobit link and fixed nuisance parameter $\lambda$. Under the flat prior $p(\beta)\propto 1$, the posterior of $\beta$ given $y_{1:n}$ (such that $\sum_{i=1}^{n}y_{i}=1$, where $\sum_{i=1}^{n}y_{i}$ is a sufficient statistic of $\beta$) is
$p(\beta\mid y_{1:n})=C\left[\frac{e^{\beta}}{\{(e^{\beta}-1)/\beta\}^{n}}\right]^{\lambda},\quad\beta\in\mathbb{R}$ (S.21)
with $C^{-1}=\int_{-\infty}^{\infty}\left[\frac{e^{\beta}}{\{(e^{\beta}-1)/\beta\}^{n}}\right]^{\lambda}\mathrm{d}\beta$, which is a proper posterior as long as $n\geq 2$. Note that for fixed $n$, the posterior distribution becomes more concentrated as $\lambda$ increases, approximately by a factor of $\lambda^{-0.5}$, since increasing $\lambda$ by a factor of 2 is equivalent to observing two independent replicates of $y_{1:n}$, based on the form of the likelihood.
Figure S.10: (Left) Examples of MCMC trace plots (only first 5,000 iterations shown) of $\beta$ for $p(\beta\mid y_{1:n})$ in (S.21) with Kolmogorov-Gamma augmentation (Right) Autocorrelation function plots with different $n$ and $\lambda$ settings.
To sample $\beta$ from $p(\beta\mid y_{1:n})$ in (S.21), we apply Algorithm 1 with Kolmogorov-Gamma augmentation, where $\lambda$ is fixed. One cycle of the MCMC algorithm from the state $(\beta^{(t)},\{\kappa_{i}^{(t)}\})$, targeting the posterior (S.21), consists of
  1. 1.
    $\kappa_{i}^{(t+1)}\mid\beta^{(t)}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}\mathrm{KG}(\lambda,\beta^{(t)})$ for $i=1,\dots,n$,
  2. 2.
    $\beta^{(t+1)}\mid\{\kappa_{i}^{(t+1)}\}_{i=1}^{n}\sim N(\lambda(1-0.5n)/K,1/K)$ where $K=\sum_{i=1}^{n}\kappa_{i}^{(t+1)}$.
For each setting of $n\in\{20,40,80\}$ and $\lambda\in\{1,10\}$, we generate $y_{1:n}$ from $Y_{i}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}\mathrm{cobin}(\beta,\lambda^{-1})$ and run the MCMC algorithm with a total of 101,000 iterations, discarding the first 1,000 samples as burn-in. Figure S.10 presents trace plots of $\beta$ (showing only the first 5,000 iterations after burn-in for clearer visualization of mixing behavior) and the corresponding autocorrelation functions (ACF). Both the trace plots and ACF indicate that the mixing of $\beta$ deteriorates as $n$ increases. For a fixed $n$, increasing $\lambda$ results in a more concentrated posterior; however, it has a negligible effect on the mixing behavior of $\beta$ as shown in the ACF.
These results suggest that in intercept-only models, when the empirical average of data $n^{-1}\sum_{i=1}^{n}y_{i}$ becomes increasingly concentrated near the boundary (which leads to more extreme intercept estimates), the Kolmogorov-Gamma augmentation scheme leads to poor mixing of $\beta$. This behavior appears to be primarily driven by the proximity of the empirical average of the data to the boundary, regardless of the empirical variance of the data that affects the nuisance parameter $\lambda$.
S.7.2 Key insights on poor mixing behavior and open problems
We illustrate that the reason for poor mixing under the Kolmogorov-Gamma augmentation with highly boundary-concentrated data is analogous to that of the Pólya-Gamma augmentation with highly imbalanced data. For the Pólya-Gamma case, the main reason of the poor mixing is that the posterior distribution of $\beta$ covers a relatively wide region, while the typical step sizes of $\beta$ induced by the Pólya-Gamma augmentation become increasingly narrower as $n$ grows (Johndrow et al., , 2019). This scale mismatch makes the typical MCMC move size of $\beta$ too small relative to the bulk of the posterior, leading to slow mixing. We show that a similar phenomenon happens for the Kolmogorov-Gamma augmentation. First, we have the following results on the posterior distribution $p(\beta\mid y_{1:n})$ (S.21) under an infinitely boundary-concentrated setting. We defer technical lemmas to the end of this section for better reading.
Theorem S.4.
Denoting $\hat{\beta}=\operatorname*{arg\,max}_{\beta}p(\beta\mid y_{1:n})$, we have $p(\hat{\beta}\mid y_{1:n})=\mathcal{O}(1/\log n)$.
Proof.
Lemma S.4 and Lemma S.5 shows that posterior density $p(\beta\mid y_{1:n})$ is essentially flat around the interval of length $\Theta(\log n)$ that contains $\hat{\beta}$. This implies $p(\hat{\beta}\mid y_{1:n})=\mathcal{O}(1/\log n)$ since the posterior density must integrate to 1. ∎
Figure S.11: Posterior densities $p(\beta\mid y_{1:n})$ in (S.21) under infinitely boundary-concentrated scenarios ($\sum_{i=1}^{n}y_{i}=1$) for $n\in\{20,40,80$} and $\lambda\in\{1,2\}$. Normalizing constants are calculated numerically.
Theorem S.4 implies that the bulk of the posterior covers the interval of length $\Omega(\log n)$. See Figure S.11 for visualization of $p(\beta\mid y_{1:n})$ with $\sum_{i=1}^{n}y_{i}=1$ when $n\in\{20,40,80\}$ and $\lambda=1$ or $\lambda=2$.
Recall that one cycle of the MCMC algorithm for the model $Y_{i}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}\mathrm{cobin}(\beta,\lambda^{-1})$ with $\sum_{i=1}^{n}y_{i}=1$ targeting the posterior (S.21), consists of steps
  1. 1.
    $\kappa_{i}^{(t+1)}\mid\beta^{(t)}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}\mathrm{KG}(\lambda,\beta^{(t)})$ for $i=1,\dots,n$,
  2. 2.
    $\beta^{(t+1)}\mid\{\kappa_{i}^{(t+1)}\}_{i=1}^{n}\sim N(\lambda(1-0.5n)/K,1/K)$ where $K=\sum_{i=1}^{n}\kappa_{i}^{(t+1)}$.
When $\beta^{(t)}$ is close to the posterior mode (e.g. $\beta^{(t)}=-n$), the mean and variance of $K=\sum_{i=1}^{n}\kappa_{i}^{(t+1)}$ satisfy $E(K\mid\beta^{(t)}=-n)=\lambda(0.5-1/n)+\mathcal{O}(ne^{-n})\approx\lambda(0.5-1/n)$ and $\operatorname*{var}(K\mid\beta^{(t)}=-n)=\lambda/(2n^{2})-2\lambda/n^{3}+\mathcal{O}(n^{-1}e^{-n})$; see Lemma S.6. Note that $\operatorname*{var}(K\mid\beta^{(t)}=-n)$ quickly shrinks to 0 as $n$ grows. Conditioning on $K=\lambda(0.5-1/n)$, we have
$\beta^{(t+1)}\mid K=\lambda(0.5-1/n)\sim N\left(-n,\frac{2}{\lambda}\frac{n}{n-2}\right).$
This means that, approximately for large $n$, the marginal transition kernel $k(\beta^{(t+1)}\mid\beta^{(t)})$ around the posterior mode ($\beta^{(t)}=-n$) involves step sizes proportional to $\lambda^{-0.5}$. This is in contrast to the fact that, as shown in Theorem S.4, the bulk of the posterior $p(\beta\mid y_{1:n})$ has increasing width with at least a $\log n$ factor. This creates a mismatch between the typical step size of the Markov chain of $\beta$ and the target distribution’s asymptotic spread. For fixed $n$, this mismatch does not appear to vanish even when the posterior becomes more concentrated with increasing $\lambda$, since larger $\lambda$ simultaneously reduces the step size at approximately the same rate, i.e., by a factor of $\lambda^{-0.5}$, as the posterior contracts. In fact, although the autocorrelation of $\beta$ remain similar, it further deteriorates the efficiency of the Markov chain, as the time required to sample a $\mathrm{KG}(\lambda,c)$ random variable based on the sum of $\lambda$ i.i.d. $\mathrm{KG}(1,c)$ variables increases linearly with $\lambda$. This phenomenon suggests several imporant and promising directions for future research, analogous to recent advances in improving data augmentation MCMC for imbalanced binary and categorical data (Duan et al., , 2018; Johndrow et al., , 2019; Zens et al., , 2024).
S.7.3 Technical lemmas
Lemma S.4.
For $n\geq 2$, $p(\beta\mid y_{1:n})$ is unimodal and attains maximum at $\hat{\beta}=-n+\mathcal{O}(1)$.
Proof.
First note that $p(\beta\mid y_{1:n})>0$ for any $\beta$, i.e. the posterior is supported on $\mathbb{R}$. By direct calculation, the derivative of $p(\beta\mid y_{1:n})$ with respect to $\beta$ satisfies $p^{\prime}(\beta\mid y_{1:n})=\lambda f(\beta)p(\beta\mid y_{1:n})$ with $f(\beta)=1-n\{1-\beta^{-1}+1/(e^{\beta}-1)\}=1-ng_{\mathrm{cobit}}^{-1}(\beta)$. Since $f^{\prime}(\beta)=-n/\beta^{2}+(ne^{\beta})/(e^{\beta}-1)^{2}<0$ for all $\beta$, $f(\beta)=0$ has the unique solution $\hat{\beta}$, which satisfies $g_{\mathrm{cobit}}^{-1}(\hat{\beta})=1/n$. Since f(n1)>0f(-n-1)>0 and f(n+2.1)<0f(-n+2.1)<0 for $n\geq 2$, $p(\beta\mid y_{1:n})$ is unimodal and $\hat{\beta}=-n+\mathcal{O}(1)$. ∎
Lemma S.5.
Let $\beta_{1}=-n+r\log(n)$ and $\beta_{2}=-n-r\log(n)$ where $r$ is some arbitrary constant satisfying $\hat{\beta}\in(\min(\beta_{1},\beta_{2}),\max(\beta_{1},\beta_{2}))$. Then $p(\beta_{1}\mid y_{1:n})/p(\beta_{2}\mid y_{1:n})\to 1$ as $n\to\infty$.
Proof.
We have
$\frac{p(\beta_{1}\mid y_{1:n})}{p(\beta_{2}\mid y_{1:n})}$ $=\left\{e^{\lambda(\beta_{1}-\beta_{2})}\left(\frac{\beta_{1}}{\beta_{2}}\right)^{\lambda n}\right\}\left\{\left(\frac{e^{\beta_{2}}-1}{e^{\beta_{1}}-1}\right)^{\lambda n}\right\}$ (S.22)
We prove the statement by showing that the two curly bracket terms above both converge to 1. For the first term, using $\beta_{1}-\beta_{2}=2r\log(n)$, and
$\log(\beta_{1}/\beta_{2})=\log(1-rn^{-1}\log n)-\log(1+rn^{-1}\log n)=-\frac{2r\log n}{n}-\mathcal{O}\left(\frac{\log^{3}n}{n^{3}}\right)$
for sufficiently large $n$ such that $|rn^{-1}\log n|<1$ using Taylor expansion $\log(1+x)=x-x^{2}/2+x^{3}/3$-$\dots$ for $|x|<1$. Thus
$\left\{e^{\beta_{1}-\beta_{2}}\left(\frac{\beta_{1}}{\beta_{2}}\right)^{n}\right\}^{\lambda}=\left\{n^{2r}\times\exp\left(n\log\frac{\beta_{1}}{\beta_{2}}\right)\right\}^{\lambda}=n^{2\lambda r}n^{-2\lambda r}\exp\left(\mathcal{O}(\frac{\log^{3}n}{n^{2}})\right)\to 1$
For the second term, using $e^{\beta_{1}}=n^{r}e^{-n}$ and $e^{\beta_{2}}=n^{-r}e^{-n}$ so that $|(e^{\beta_{1}}-e^{\beta_{2}})/(1-e^{\beta_{1}})|=e^{-n}|n^{r}-n^{-r}|/|1-e^{\beta_{1}}|=\mathcal{O}(e^{-n}n^{|r|})$. Since
$\left(\frac{e^{\beta_{2}}-1}{e^{\beta_{1}}-1}\right)^{n}=\exp\left(n\log\left(\frac{e^{\beta_{2}}-1}{e^{\beta_{1}}-1}\right)\right)=\exp\left(n\log\left(1+\frac{e^{\beta_{1}}-e^{\beta_{2}}}{1-e^{\beta_{1}}}\right)\right),$
using $|\log(1+x)|\leq 2|x|$ for $|x|<1/2$, we have $n|\log\left(1+\frac{e^{\beta_{1}}-e^{\beta_{2}}}{1-e^{\beta_{1}}}\right)|\leq 2n|\frac{e^{\beta_{1}}-e^{\beta_{2}}}{1-e^{\beta_{1}}}|\to 0$, which implies $\left(\frac{e^{\beta_{2}}-1}{e^{\beta_{1}}-1}\right)^{n}\to 1$, thus $\left(\frac{e^{\beta_{2}}-1}{e^{\beta_{1}}-1}\right)^{\lambda n}\to 1$. ∎
Lemma S.6.
If $\kappa\sim\mathrm{KG}(b,c)$ and $c\neq 0$, we have
$E(\kappa)$ $=bc^{-2}\{(c/2)\coth(c/2)-1\}$
$\operatorname*{var}(\kappa)$ $=\frac{b}{4c^{4}}\mathrm{csch}^{2}(c/2)(c^{2}+c\sinh(c)-4\cosh(c)+4)$
If $c=0$, $E(\kappa)=b/12$ and $\operatorname*{var}(\kappa)=b/360$. For large $c>0$, recalling $\mathrm{KG}(b,c)\stackrel{{\scriptstyle d}}{{=}}\mathrm{KG}(b,-c)$,
$E(\kappa)$ $=b(0.5c^{-1}-c^{-2})+\mathcal{O}(c^{-1}e^{-c})$
$\operatorname*{var}(\kappa)$ $=b(0.5c^{-3}-2c^{-4})+\mathcal{O}(c^{-2}e^{-c})$
Proof.
These are derived straightforwardly from the definition of the Kolmogorov-Gamma as an infinite convolution of gammas, using $\sum_{k=1}^{\infty}1/\{k^{2}+c^{2}/(4\pi^{2})\}=(\pi^{2}/c)\coth(c/2)-2\pi^{2}/c^{2}$ (Gradshteyn and Ryzhik, , 2014, §1.421.4). Asymptotic derivations are straightforwardly follow from $(c/2)\coth(c/2)=c/2+ce^{-c}/(1-e^{-c})$, $\mathrm{csch}^{2}(c/2)=4e^{c}/(e^{c}-1)^{2}=4e^{-c}(1+2e^{-c}+3e^{-2c}+\cdots)$. ∎